1. The business offers catering services. For the period ended, it collected 250,000 cash and still has receivables amounting to 80,000 for the total services rendered. How will you compute the revenue of the business for the period? How much is its income?
2. The assets of the business compromise of 200,000 cash and 1,800,000 Land and Building. The only obligation by the business is its loan to Rural Bank of Baao related to its acquisition of land and building. The owner's equity is 1,300,000. How much is the liability of the business? How did you come up with the amount?
1
Expert's answer
2021-09-14T07:30:02-0400

Solution:

1.). The revenue of the business for the period will be calculated by adding the total cash collected and the account receivables:

Total revenue = Total cash collected + account receivable

Total revenue = 250,000 + 80,000 = 330,000

Total revenue for the period = 330,000

 

The income of the business is the total amount collected as at the end of the period.

Amount of income collected = 250,000

Income of the business = 250,000

 

2.). Total Assets = Total liabilities + Owners Equity

The total liabilities (Loan to Rural Bank) = 700,000

This is derived as follows:

Total Assets = Cash + Land & building = 200,000 + 1,800,000 = 2,000,000

Owners Equity = 1,300,000

Total liabilities = Total Assets – Owners Equity = 2,000,000 – 1,300,000 = 700,000

The total liabilities (Loan to Rural Bank of Baao) = 700,000